OpenStudy (anonymous):

How can i solve 3n - 4 = 11

OpenStudy (anonymous):

3n-4=11 +4 +4 3n=15 /3 /3 n=5

OpenStudy (anonymous):

3n-4=11 +4 +4 (make 4 positive and move it to the other side) 3n=15 /3 /3 (take 3 and divide it with 15) end up with n=5

OpenStudy (anonymous):

Thank you so much

OpenStudy (anonymous):

np